# FactSelfCheck: Fact-Level Black-Box Hallucination Detection for LLMs

### v1

@@ Appendix C Passage-level Score: Size of Knowledge Graphs Extracted from Samples
$h\in\mathcal{E}_{S}\land r\in\mathcal{R}_{S}\land t\in\mathcal{E}_{S}\}$
The passage-level score is then calculated as:
$\mathcal{H}_{\text{passage}}(p)=\frac{1}{|S|}\sum_{s\in S}-|KG_{s}^{\prime}|$ (14)
The rationale behind this score is that the size of $KG_{s}^{\prime}$ reflects how well the samples align with the original response $p$. A smaller $KG_{s}^{\prime}$ indicates that the sample contains facts than those in the original response, suggesting potential hallucinations. In distinction from the previously defined scores, this metric operates outside the standard $[0,1]$ range.
Results
The method achieved Pearson and Spearman correlation scores of $72.35$ and $71.18$, respectively. When compared to the main results in Table 2, this approach shows lower correlation scores than LLM-based methods but outperforms the frequency-based method. It is important to note that while it is the passage-level score, frequency-based methods operate at the fact level. Nevertheless, this approach offers the most computationally efficient way to calculate passage-level scores among tested methods utilizing knowledge graphs, as it does not require any LLM calls during factual consistency assessment.
Appendix D Case Study of FactSelfCheck vs SelfCheckGPT
@@
Table 5 and Figure 5 present a comparative case study of predictions made by FactSelfCheck and SelfCheckGPT. The table contains an external Wikipedia biography, sentences from the response, facts extracted from the response, and the predictions of both methods. This comparison demonstrates that fact-level detection provides more detailed information about the factuality of the response. We observe that LLM did not hallucinate all facts, as some are consistent with the external Wikipedia biography. However, when using sentence-level detection, we cannot distinguish between correct and hallucinated facts – all sentences are predicted as hallucinated.
Appendix E SelfCheckGPT with Enhanced Prompt
@@
To ensure a fair comparison between FactSelfCheck and SelfCheckGPT, we conducted an additional experiment using an enhanced prompt for SelfCheckGPT. The original SelfCheckGPT prompt is relatively simple, while our FactSelfCheck prompts are more elaborate, directly allowing reasoning and inference of new facts, and providing examples. These characteristics could potentially increase the performance of methods. We designed an alternative prompt for SelfCheckGPT, presented in Prompt 12, that incorporates these features, making it similar to our FactSelfCheck prompts.
@@
For sentence-level detection, the enhanced prompt for SelfCheckGPT achieved an AUC-PR score of $93.38$, slightly lower than the original prompt’s $93.60$. This minimal difference indicates that SelfCheckGPT’s performance is not significantly affected by prompt design in our experimental setting. In fact, the enhanced prompt even lowered the performance rather than increasing it, despite its more sophisticated design. For passage-level detection, we observed similar results, with the enhanced prompt achieving Pearson and Spearman correlation scores of $82.36$ and $76.71$, respectively, compared to the original prompt’s $83.12$ and $78.65$.
@@
Due to these findings, we chose to use the original prompt for SelfCheckGPT in our main experiments. These consistent results confirm that our comparison between methods is fair, as the enhanced prompt did not improve the performance of SelfCheckGPT.
Appendix F FactSelfCheck Prompts
In this section, we present the prompts used in FactSelfCheck. These include prompts for the entity extraction (Prompt 1), the relation extraction (Prompt 2), and the knowledge graph extraction from a sentence (Prompt 3) or a sample (Prompt 4). Additionally, we provide prompts for assessing factual consistency in FactSelfCheck-KG (Prompt 5) and FactSelfCheck-Text (Prompt 6).
Appendix G Correction Prompts
The following prompts are instructions for correcting hallucinated responses generated by the LLM. The prompts are for the baseline (Prompt 7), sentence-level (Prompt 8), and fact-level (Prompt 9) correction. The prompts take several inputs: {generated_sentences} containing the response as a numbered list of sentences, {incorrect_sentences} listing the hallucinated sentences with their corresponding numbers from the generated text, and {incorrect_facts} providing a formatted list of hallucinated facts organized by sentence. The {format} field specifies the output format as an empty numbered list with a length equal to the number of sentences in the {generated_sentences} field.
@@ Appendix G Correction Prompts
Table 5: Comparison of fact-level FactSelfCheck with sentence-level SelfCheckGPT. An external Wikipedia biography is provided to analyse the correctness of the methods. The red value indicates hallucinations, and the green value indicates factual correctness. The facts were classified using a threshold of $0.4$ utilizing FactSelfCheck, and the sentences were classified using a threshold of $0.75$ with SelfCheckGPT. These thresholds achieved the highest F1-scores in fact-level (Appendix B) and sentence-level (Section 5.1) evaluation, respectively.

### v2

@@ D.1 Sentence-Level Detection
$\mathcal{H}$) - 79.02 49.18 64.10
[…]
AttentionScore (Absolute) - 77.95 42.23 60.09
RandomSentence - 72.96 27.04 50.00
Fact-level methods (ours)
FSC-Text max 92.45 65.55 79.00
FSC-KG (LLM-based) max 91.82 64.64 78.23
FSC-Text mean 91.01 63.77 77.39
FSC-KG (LLM-based) mean 90.24 62.95 76.60
FSC-KG (Frequency-based) max 88.48 53.86 71.17
FSC-KG (Frequency-based) mean 88.25 55.27 71.76
RandomFact mean 74.22 29.74 51.98
Table 6: WikiBio: Extended results on the sentence-level hallucination detection task. Comparison of sentence-level and fact-level methods based on AUC-PR scores for Hallucination, Factuality, and their Average. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores. The results are sorted by AUC-PR scores for Hallucination.
Method Agg. AUC-PR
Hallucination Factuality Avg.
Sentence-level methods
SCGPT (Prompt) - 46.91 88.39 67.65
SCGPT (NLI) - 32.58 85.64 59.11
$\mathcal{H}$) - 28.22 81.90 55.06
[…]
AttentionScore (Relative) - 24.17 82.12 53.15
$\mathcal{H}$) - 23.80 80.62 52.21
[…]
AttentionScore (Absolute) - 22.19 81.71 51.95
RandomSentence - 21.70 78.30 50.00
Fact-level methods (ours)
FSC-KG (Frequency-based) max 48.52 80.21 64.36
FSC-Text max 42.80 86.15 64.47
FSC-KG (LLM-based) max 40.63 85.01 62.82
FSC-Text mean 37.13 86.09 61.61
FSC-KG (Frequency-based) mean 36.16 79.92 58.04
FSC-KG (LLM-based) mean 35.81 84.65 60.23
RandomFact mean 21.22 77.77 49.50
Table 7: FavaMultiSamples: Extended results on the sentence-level hallucination detection task. Comparison of sentence-level and fact-level methods based on AUC-PR scores for Hallucination, Factuality, and their Average. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores. The results are sorted by AUC-PR scores for Hallucination.
Figure 2: WikiBio: Precision-recall curve for the sentence-level hallucination and factuality detection.
Figure 3: FavaMultiSamples: Precision-recall curve for the sentence-level hallucination and factuality detection.
D.2 Effect of Sample Size on Detection Performance
We investigated the impact of varying the number of samples on the detection performance of our method. Specifically, we evaluated the performance at the sentence level by changing the number of samples from $1$ to $20$. We compared our method with SelfCheckGPT.
Figure 4: WikiBio: Impact of sample size on both hallucination and factuality detection performance for different methods.
Figure 5: FavaMultiSamples: Impact of sample size on both hallucination and factuality detection performance for different methods.
Figure 4 illustrates that that on WikiBio, FactSelfCheck exhibits similar behavior to SelfCheckGPT regarding sample requirements. The performance improves dramatically with up to $5$ samples, after which the improvement curve flattens. While additional samples continue to yield benefits, these improvements become incremental, with modest gains observed up to $20$ samples. Factuality detection exhibits similar patterns to hallucination detection. This pattern confirms that more samples provide better evidence for accurate detection across both metrics.
Figure 5 shows a different pattern for FavaMultiSamples, where hallucination detection performance decreases with more samples. This occurs because methods with few samples tend to overestimate hallucination scores, interpreting normal variations as potential hallucinations. As sample size increases, methods become more conservative in their scoring, leading to better calibration. This pattern likely stems from FavaMultiSamples having shorter sentences with fewer facts compared to WikiBio (see Section D.3).
The distinct patterns between datasets highlight that sampling effects are context-dependent. While more samples universally improve calibration quality, the impact on raw performance metrics depends on dataset characteristics and initial score distributions. It’s important to note that AUC-PR, while useful, has limitations. A classifier that consistently returns the same score can achieve a high AUC-PR value, which may not reflect true discriminative ability. Therefore, the decrease in AUC-PR with more samples might actually indicate better calibration and more meaningful score distributions, rather than worse performance.
D.3 Evaluation of Intermediate Steps
Entities/ Relations/ Facts/
passage passage sentence
WikiBio
mean 24.85 21.18 3.24
min 11 5 0
max 282 163 85
# 0 el. 0 0 22
% 0 el. 0.00% 0.00% 1.15%
FavaMultiSamples
mean 32.08 50.34 2.97
min 1 0 0
max 312 237 102
# 0 el. 0 2 344
% 0 el. 0.00% 0.43% 6.08%
Table 8: Statistics of intermediate steps in FactSelfCheck across both datasets. The table shows the distribution of entities per passage, relations per passage, and facts per sentence, including occurrences of entries with 0 elements.
Our method consists of multiple steps that cannot be directly evaluated due to the lack of human annotations. While previous sections evaluated the complete pipeline, to strengthen our study we also analyzed and validated statistics from intermediate steps. We examined the number of entities and relations per passage, along with facts per sentence, calculating mean, minimum, maximum values, and the count and percentage of entries with zero elements. Sentences with no facts are particularly important as FactSelfCheck assigns them a default score of $0.5$ (see Section 4.3).
Table 8 reveals that both datasets have high mean numbers of entities and relations per passage, indicating that knowledge graph construction is not constrained by earlier steps. While WikiBio shows acceptable minimum values for entities and relations per passage, FavaMultiSamples exhibits notably lower minimums that could impact knowledge graph extraction performance. The percentage of sentences without extracted facts is relatively low in WikiBio ($1.15\%$) but more substantial in FavaMultiSamples ($6.08\%$), potentially affecting detection accuracy. Both datasets show considerable variability, with some passages containing over 200 entities and relations, highlighting the diverse complexity of the analyzed generated responses.
Appendix E Computional complexity
Method Number of LLM calls
FSC-KG (Freq.) $2+|U|+|S|$
FSC-KG (LLM) $2+|U|+|S|+|KG_{p}|\times|S|$
FSC-Text $2+|U|+|KG_{p}|\times|S|$
SCGPT-Prompt $|U|\times|S|$
Table 9: Computational complexity of different methods in terms of number of LLM calls required. $U$ is the set of sentences in the generated passage, $S$ is the set of stochastic LLM response samples, and $KG_{p}$ is the knowledge graph containing all extracted facts.
While FactSelfCheck is more granular than SelfCheckGPT, this comes with increased computational costs. In Table 9 we compare all variants of FactSelfCheck with SelfCheckGPT (Prompt) in terms of the number of LLM calls required. Every variant of FactSelfCheck requires additional calls for entity and relation extraction (constant $2$), followed by knowledge graph construction for each sentence ($|U|$), then it assesses factual consistency of each fact in different ways with varying complexity. As noted in Limitations, we did not optimize for token usage, and future work could merge steps to reduce complexity while maintaining the fine-grained insights our method provides.
Appendix F Case Study of FactSelfCheck vs SelfCheckGPT
@@
Table 10 and Figure 6 present a comparative case study of predictions from WikiBio made by FactSelfCheck and SelfCheckGPT. The table contains an external Wikipedia biography, sentences from the response, facts extracted from the response, and the predictions of both methods. This comparison demonstrates that fact-level detection provides more detailed information about the factuality of the response. We observe that LLM did not hallucinate all facts, as some are consistent with the external Wikipedia biography. However, when using sentence-level detection, we cannot distinguish between correct and hallucinated facts – all sentences are predicted as hallucinated.
Appendix G SelfCheckGPT with Enhanced Prompt
@@
To ensure a fair comparison between FactSelfCheck and SelfCheckGPT, during preliminary studies, we conducted an additional experiment using an enhanced prompt for SelfCheckGPT. The original SelfCheckGPT prompt is relatively simple, while our FactSelfCheck prompts are more elaborate, directly allowing reasoning and inference of new facts, and providing examples. These characteristics could potentially increase the performance of methods. We designed an alternative prompt for SelfCheckGPT, that incorporates these features, making it similar to our FactSelfCheck prompts.
@@
For sentence-level detection on WikiBio, the enhanced prompt for SelfCheckGPT achieved an AUC-PR score of $93.38$, slightly lower than the original prompt’s $93.60$. This minimal difference indicates that SelfCheckGPT’s performance is not significantly affected by prompt design in our experimental setting. In fact, the enhanced prompt even lowered the performance rather than increasing it, despite its more sophisticated design.
@@
Due to these findings, we chose to use the original prompt for SelfCheckGPT in our experiments. These consistent results confirm that our comparison between methods is fair, as the enhanced prompt did not improve the performance of SelfCheckGPT.
@@
Table 10: Comparison of fact-level FactSelfCheck with sentence-level SelfCheckGPT. An external Wikipedia biography is provided to analyse the correctness of the methods. The red value indicates hallucinations, and the green value indicates factual correctness. The facts were classified using a threshold of $0.4$ utilizing FactSelfCheck, and the sentences were classified using a threshold of $0.75$ with SelfCheckGPT. These thresholds achieved the highest F1-scores in fact-level (Appendix C) and sentence-level (Section 5.1) evaluation, respectively.